Alex is 16 year old and his sister is 5 years old. In how many years will he be exactly two times as old as his sister?
Answer by checkley77(12844)

16+x=2(5+x)
16+x=10+2x
x-2x=10-16
-x=-6
x=6 years from now Alex will be 2 times as old as his sister.
Proof:
16+6=2(5=6)
22=2*11
22=22